Nymonka09744:

Tyra has 1223 cups of sugar. If r is the amount of sugar called for in one batch of cookies, the expression 1223 ÷ r can be used to find how many batches Tyra can make with the sugar she has. If r is 23 cup, how many batches of cookies can Tyra make?

Joe348:

\[12 \frac{ 2 }{ 3 } \div \frac{ 2 }{ 3 }=19\] She can make 19 batches!

Nymonka09744:

@joe348 wrote:
\[12 \frac{ 2 }{ 3 } \div \frac{ 2 }{ 3 }=19\] She can make 19 batches!
thanks alot i got a 40% because of u

Joe348:

There's no possible way I got you an 40%

Joe348:

I'm sure you had other questions to and that's on you not me, I helped with the question you asked that's it